# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v3

@@ 2.2. The Duals of Quasi-Cyclic Codes
[HKS, Proposition 7.3.5] Let $\mathcal{C}$ be a QC code with CRT decomposition as (10). Then its dual code $\mathcal{C}^{\perp}$ (under the Euclidean inner product) is of the form
(21) $\mathcal{C}^{\perp_{E}}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{{}^{\prime}\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}^{\perp_{H}}\right)$
Since the duality of $\mathcal{C}$ can be determined from its constituent codes, the following results provide insight to extended constructions of self-orthogonal codes.
@@
[HillNewton, Lemma 2.10] Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ code over $\mathbb{F}_{q}$ .
In particular, a version of Lemma 2.7 regarding self-orthogonal codes is presented below.
@@
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code $mC$ as $m$-copies of $C$ which will play an important role in the next section.
2.3. A Lower Bound of the Minimum Distance of a QC Code
Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of a $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By [cem, Lemma 4.1], the trace-formula described in (13) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, looking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of a $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
$BZ(D^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i=1,2,...,h\}.$
In [BastosAlvarezFloresSalerno], the authors called $D$ the cyclic code associated to $\mathcal{C}$. Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
$1\leq i_{1}<i_{2}<\cdots<i_{t}\leq h.$
Let $D_{I}=D_{i_{1},...,i_{t}}\subset D$ be the cyclic subcode of $D$ whose dual’s basic zero set is
(22) $BZ(D_{I}^{\perp})=\{\alpha^{-u_{a_{i}}}\mid i\in I\}.$
Define the following quantity:
(23) $R_{I}(\mathcal{C})$ $=R_{i_{1},...,i_{t}}(\mathcal{C})$
(24) $:=\begin{cases}d(C_{i_{1}})d(D_{i_{1}})&\text{if $t=1$}\\ \\ \left(d(C_{i_{1}})-d(C_{i_{2}})\right)d(D_{i_{1}})+\left(d(C_{i_{2}})-d(C_{i_{3}})\right)d(D_{i_{1},i_{2}})+\\ \hskip 128.0374pt\vdots&\text{if $t\geq 2$}\\ +\left(d(C_{i_{t-1}})-d(C_{i_{t}})\right)d(D_{i_{1},i_{2},...i_{t-1}})+d(C_{i_{t}})d(D_{i_{1},i_{2},...,i_{t}}).\end{cases}$
The next result uses the $R_{I}$ to provide a lower bound to the minimum distance of a $\ell$-QC code $\mathcal{C}$ based on the minimum distances of its constituent codes and the cyclic code (and its cyclic subcodes) associated to $\mathcal{C}$. This lower bound is usually referred as GO-bound where relevant details can be checked in [cem]. First, we assume that the minimum distances of the nonzero constituent codes $C_{1},C_{2},...,C_{h}$ of $\mathcal{C}$ satisfy the following condition
(25) $d(C_{1})\geq d(C_{2})\geq...\geq d(C_{h}).$
Next, we use the slightly modified version of the original theorem which will be useful in determining the minimum distances of self-orthogonal (resp. dual containing or self dual) QC codes obtained from our construction in Section 3.
Theorem 2.9 (Güneri-Özbudak bound).
[cem, Theorem 4.8] Let $\mathcal{C}$ be a QC code as in (10) with the assumption (25). Then we have
(26) $d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{R_{h}(\mathcal{C}),R_{h-1,h}(\mathcal{C}),R_{h-2,h-1,h}(\mathcal{C}),...,R_{1,2,...,h}(\mathcal{C})\}.$
Remark 2.
  1. (i)
    If the number of nonzero constituent codes is two, the GO-bound (26) becomes
    $d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{d(C_{2})d(D_{2}),\left(d(C_{1})-d(C_{2})\right)d(D_{1})+d(C_{2})d(D)\},$
    where $D=D_{1,2}$.
  2. (i)
    If the number of nonzero constituent codes is three, it becomes
    $d(\mathcal{C})\geq d_{\text{GO}}(\mathcal{C}):=\min\{R_{3},R_{2,3},R_{1,2,3}\},$
    where
    $R_{3}(\mathcal{C})$ $=$ $d(C_{3})d(D_{3})$
    $R_{2,3}(\mathcal{C})$ $=$ $(d(C_{2})-d(C_{3}))d(D_{2})+d(C_{3})d(D_{2,3})$
    $R_{1,2,3}(\mathcal{C})$ $=$ $(d(C_{1})-d(C_{2}))d(D_{1})+(d(C_{2})-d(C_{3}))d(D_{1,3})+d(C_{3})d(D),$
    and $D=D_{1,2,3}$.
2.4. Constructions of Quantum Error-Correcting Codes from Classical Ones
From [Ketkar-Klappenecker-Kumar-Sarvepalli2006], the use of classical codes has became one of the standard methods for constructing quantum error-correcting codes. A quantum error-correcting code (QECC) $Q$ is a $K$-dimensional subspace of $(\mathbb{C}^{q})^{\otimes n}$. If $Q$ has minimum distance $d$, then we say that $Q$ is an $n,K,d]]_{q}$ code. If $K=q^{k}$, we write $n,k,d]]_{q}$. The length $n$, the dimension $K$, and minimum distance $d$ are the parameters of $Q$.
If $Q$ is a quantum code with minimum distance $d$, $Q$ can detect up to d1d-1 quantum errors, and can correct up to $\lfloor\frac{d-1}{1}\rfloor$ quantum errors. The celebrated Singleton bound [LaGuardia2021, Lemma 3.5.1] says that the parameters of an $n,K,d]]_{q}$ quantum code satisfy
(27) $\log_{q}(K)\leq n-2d+2$
@@ 2.4. Constructions of Quantum Error-Correcting Codes from Classical Ones
This paper focuses on the construction and existence of quantum stabilizer codes, where a stabilizer (quantum) code $Q\neq\{\textbf{0}\}$ is the common eigenspace of a commutative subgroup of the error group generated by a nice basis of $(\mathbb{C}^{q})^{\otimes n}$. The code $Q$ is said to be pure to d, or simply pure, if and only if its stabilizer group does not contain non-scalar matrices of weight less than its minimum distance $d$. (See Definitions 3.5.7 and 3.5.8 in [LaGuardia2021].) Moreover, a code $Q$ is said to be impure if and only if there are non-identity stabilizer elements of weight less than the minimum distance. In 1996, the following construction was introduced by Calderbank and Shor [Calderbank-Shor1996], and Steane [Steane1996], and is notably the most direct link between classical and quantum coding theory.
Theorem 2.10 (CSS construction [Ketkar-Klappenecker-Kumar-Sarvepalli2006]).
@@
In the special case when $C_{2}=C_{1}$, we have that $C_{1}$ is a dual-containing code. This yields
@@
[Aly-Klappenecker-Sarvepalli2007, Ketkar-Klappenecker-Kumar-Sarvepalli2006] Let $C$ be a classical $[n,k,d]$ code over $\mathbb{F}_{q}$. If $C$ is dual-containing, then there exists a quantum stabilizer code $Q$ with parameters $n,2k-n,\geq d]]_{q}$ that is pure to $d$. Moreover, if the minimum distance of $C^{\perp}$ exceeds $d$, then the quantum code $Q$ has minimum distance $d$.
In recent years, researchers in coding theory have worked diligently to construct good (e.g., quantum MDS) quantum codes from classical Euclidean, Hermitian, or symplectic self-orthogonal, dual-containing, or self-dual codes [Abdukhalikov-Bag-Panario2023, Ball21, Cao24, DinhLeNguyenTansuchat21, FangFu19, GalindoHernando, galindo, GaoSunYanZhao25, JinXing13, LiuLiu21, ShiYueChang18, TianZhu19, WanZhengZhu24]. Moreover, the following results (which will be used in examples in Section 4) present new quantum code constructions utilizing the lengthening, shortening, and dimension reduction of a given stabilizer quantum code.
@@
[Ketkar-Klappenecker-Kumar-Sarvepalli2006, Lemmas 69, 70, 71, and Corollary 73]
  1. (1)
@@
  4. (4)
    Suppose that an $n,q^{k},d]]_{q}$ and an $n^{\prime},q^{k^{\prime}},d^{\prime}]]$ stabilizer code exist. Then there exists an $n+n^{\prime},q^{k+k^{\prime}},\min\{d,d^{\prime}\}]]_{q}$ stabilizer code.
3. A Construction of ESO/EDC/ESD QC codes from their constituent codes
In this section, we discuss the main results of this paper where we construct infinite families $\{\mathcal{C}_{u}\}_{u\geq 1}$ of QC codes that are self-orthogonal (or self-dual, or dual-containing) with respect to the Euclidean inner product. Using the notion of Galois closed codes, we also propose a construction of QC codes that are self-orthogonal (or self-dual, or dual-containing) with respect to the Hermitian inner product. Such constructions of QC codes will be utilized to show the existence of QECCs with good parameters in Section 4.
3.1. Constructions of Infinite Families of QC Codes
We propose a recursive construction of infinite families $\{\mathcal{C}_{u}\}_{u\geq 1}$ of ESO QC codes using its constituent codes. To provide a careful description of our construction, we first describe a general case of the construction and then specify certain conditions on the constituent codes to explicitly describe the code dimension and lower bound of its minimum distance.
Let $x^{m}-1$ factor into $2t+s$ irreducible factors, namely
(28) $x^{m}-1=\delta g_{1}g_{1}^{*}\cdots g_{t}g_{t}^{*}f_{1}\cdots f_{s-1}f_{s},$
where $\delta\in\mathbb{F}_{q}$ is nonzero, $g_{1},g_{1}^{*},...,g_{t},g_{t}^{*}$ are the factors grouped in pairs with their reciprocal polynomials for each $1\leq j\leq t$, and $f_{1},...,f_{s}$ are self-reciprocal polynomials for $1\leq i\leq s$, where $f_{s}=x-1$. Then,
(29) $R^{\ell}_{q,m}\cong\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{1}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{1}^{*}\rangle}\right)^{\ell}\oplus\cdots\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{t}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{t}^{*}\rangle}\right)^{\ell}\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{1}\rangle}\right)^{\ell}\oplus\cdots\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{s-1}\rangle}\right)^{\ell}\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}.$
3.1.1. Recursive Construction:
Let $\mathcal{C}_{1}\subseteq\mathbb{F}_{q}^{m\ell}$ be an $\ell$-QC whose first $2t$ constituent codes are direct sums of arbitrary codes and their duals, i.e.,
(30) $\bigoplus_{j=1}^{t}\left(C^{\prime}_{1j}\oplus{C^{\prime}}_{1j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}$
@@ 3.1.1. Recursive Construction:
(31) $\left(\bigoplus_{i=1}^{s-1}C_{1i}\right)\oplus C_{1s}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$
where the codes $C_{1i}$ are HSO codes for $1\leq i\leq s-1$ and $C_{1s}$ is ESO (see Remark 1). Thus,
(32) $\mathcal{C}_{1}\cong\bigoplus_{j=1}^{t}\left(C^{\prime}_{1j}\oplus{C^{\prime}}_{1j}^{\perp_{E}}\right)\oplus\left(\bigoplus_{i=1}^{s-1}C_{1i}\right)\oplus C_{1s}\subset R_{q,m}^{\ell}.$
We also assume that the minimum distances of these constituent codes satisfy the inequalities in (25), i.e.,
(33) $d(C_{11}^{\prime})\geq d((C_{11}^{\prime})^{\perp_{E}})\geq...\geq d(C_{1t}^{\prime})\geq d((C_{1t}^{\prime})^{\perp_{E}})\geq d(C_{11})\geq...\geq d(C_{1,s-1})\geq d(C_{1s}).$
As the $C_{1i}$ are HSO for each $1\leq i\leq s-1$ and $C_{1s}$ is ESO, Theorem 2.5 tells us that $\mathcal{C}_{1}$ is an ESO $\ell$-QC code of length $m\ell$. Now using $\mathcal{C}_{1}$, we construct another QC code $\mathcal{C}_{2}$ as follows: Let
(34) $R^{m\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}.$
Let $\mathcal{C}_{2}\subseteq\mathbb{F}_{q}^{m^{2}\ell}$ be the $\ell$-QC code whose first $2t$ constituent codes are the following direct sum of arbitrary codes and their duals,
(35) $\bigoplus_{j=1}^{t}\left(C^{\prime}_{2j}\oplus{C^{\prime}}_{2j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell},$
@@
(36) $\left(\bigoplus_{i=1}^{s-1}C_{2i}\right)\oplus\mathcal{C}_{1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}$
@@
where the $C_{2i}$ are HSO codes for $1\leq i\leq s-1$ and $\mathcal{C}_{1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}\cong\mathbb{F}_{q}^{m\ell}$ is an ESO constituent code of $\mathcal{C}_{2}$. Thus,
@@
(37) $\mathcal{C}_{2}\cong\bigoplus_{j=1}^{t}\left(C^{\prime}_{2j}\oplus{C^{\prime}}_{2j}^{\perp_{E}}\right)\oplus\left(\bigoplus_{i=1}^{s-1}\;C_{2i}\right)\oplus\mathcal{C}_{1}$
@@
Again by Theorem 2.5, $\mathcal{C}_{2}$ is an ESO $\ell$-QC code of length $m^{2}\ell$. To extend this recursive construction, we have that any $u\geq 2$,
@@
(38) $R^{m^{u-1}\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}.$
@@
and define an $\ell$-QC code $\mathcal{C}_{u}$ in $\mathbb{F}_{q}^{m^{u}\ell}$ as
(39) $\mathcal{C}_{u}\cong\left(C^{\prime}_{u1}\oplus{C^{\prime}}_{u1}^{\perp}\right)\oplus\left(C^{\prime}_{u2}\oplus{C^{\prime}}_{u2}^{\perp}\right)\oplus\cdots\oplus\left(C^{\prime}_{u,t}\oplus{C^{\prime}}_{u,t}^{\perp}\right)\oplus C_{u1}\oplus\cdots\oplus C_{u,s-1}\oplus\mathcal{C}_{u-1},$
where for each $u$, the first $2t$ arbitrary constituent codes are

### v4

@@ 2.3 The Duals of Quasi-Cyclic Codes
(HKS, Proposition 7.3.5) Let $\mathcal{C}$ be a QC code with CRT decomposition as (9). Then its Euclidean dual $\mathcal{C}^{\perp_{E}}$ is of the form
$\mathcal{C}^{\perp_{E}}\cong\left(\bigoplus_{j=1}^{t}(C_{j}^{\prime}\oplus C_{j}^{{}^{\prime}\perp_{E}})\right)\oplus\left(\bigoplus_{i=1}^{s}C_{i}^{\perp_{H}}\right).$
Since the duality of $\mathcal{C}$ can be determined from its constituent codes, we will use the following results in Li for constructing self-orthogonal codes.
lightgray Thus, when we apply Proposition 2.10 to Jensen’s concatenated description in (17), we obtain
$\mathcal{C}^{\perp_{E}}\cong\left(\bigoplus_{j=1}^{t}\left(D_{j}^{\prime}\square C_{j}^{\prime}\oplus D_{j}^{{}^{\prime}\perp_{E}}\square C_{j}^{{}^{\prime}\perp_{E}}\right)\right)\oplus\left(\bigoplus_{i=1}^{s}D_{i}\square C_{i}^{\perp_{H}}\right).$ (21)
Such representation of a QC code $\mathcal{C}$ is vital when dealing with the duality of quasi-cyclic codes $\mathcal{C}$.
@@
(HillNewton, Lemma 2.10) Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ code over $\mathbb{F}_{q}$.
@@
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code $mC$ as $m$-copies of $C$ which will play an important role in the next section when determining the dimension of the QC codes in our infinite family.
2.4 Constructions of Quantum Error-Correcting Codes from Classical Ones
From Ketkar-Klappenecker-Kumar-Sarvepalli2006, the use of classical codes has became one of the standard methods for constructing quantum error-correcting codes. A quantum error-correcting code (QECC) $Q$ is a $K$-dimensional subspace of $(\mathbb{C}^{q})^{\otimes n}$. If $Q$ has minimum distance $d$, then we say that $Q$ is an $n,K,d]]_{q}$ code. If $K=q^{k}$, we write $n,k,d]]_{q}$. The length $n$, the dimension $K$, and minimum distance $d$ are the parameters of $Q$.
f $Q$ is a quantum code with minimum distance $d$, $Q$ can detect up to d1d-1 quantum errors, and can correct up to $\lfloor\frac{d-1}{2}\rfloor$ quantum errors. The celebrated Singleton bound (LaGuardia2021, Lemma 3.5.1) says that the parameters of an $n,K,d]]_{q}$ quantum code satisfy
$\log_{q}(K)\leq n-2d+2$
@@ 2.4 Constructions of Quantum Error-Correcting Codes from Classical Ones
This paper focuses on the existence of quantum stabilizer codes, where a stabilizer (quantum) code $Q\neq\{\textbf{0}\}$ is the common eigenspace of a commutative subgroup of the error group generated by a nice basis of $(\mathbb{C}^{q})^{\otimes n}$. The code $Q$ is said to be pure to d, or simply pure, if and only if its stabilizer group does not contain non-scalar matrices of weight less than its minimum distance $d$. (See Definitions 3.5.7 and 3.5.8 in LaGuardia2021.) Moreover, a code $Q$ is said to be impure if and only if there are non-identity stabilizer elements of weight less than the minimum distance. n other words,
[…]
Definition 2.13.
The code $Q$ is said to be pure to l if and only if its stabilizer group does not contain non-scalar matrices of weight less than $l$; $Q$ is pure if and only if its pure to its minimum distance. G.: I did not understand this definition :/
Lemma 2.14.
(Ketkar-Klappenecker-Kumar-Sarvepalli2006, Corollary 80) The parameters of an $n,K,d]]_{q}$ stabilizer code with $K>1$ satisfy the inequality
$K\leq q^{n-2d+2}.$ (23)
In 1996, the following construction was introduced by Calderbank and Shor Calderbank-Shor1996, and Steane Steane1996, and is notably the most direct link between classical and quantum coding theory.
Theorem 2.15 (CSS construction Ketkar-Klappenecker-Kumar-Sarvepalli2006).
@@
theorem[CSS construction] Let $C_{1}$ and $C_{2}$ be classic linear codes over $\mathbb{F}_{q}$. If $C_{2}\subset C_{1}\subset\mathbb{F}_{q}^{n}$, there exisst a stabilizer quantum code with parameters
$n,\dim(C_{1})-\dim(C_{2}),\min\{d(C_{1}\backslash C_{2}),d(C_{2}^{\perp}\backslash C_{1}^{\perp})\}]]_{q}$ (24)
where $C_{i}^{\perp}$ is the Euclidean dual of $C_{i}$, for $i=1,2$. In the special case when $C_{2}=C_{1}$, we have that $C_{1}$ is a dual-containing code. This yields the following result.
@@
Aly-Klappenecker-Sarvepalli2007; Ketkar-Klappenecker-Kumar-Sarvepalli2006 Let $C$ be a classical $[n,k,d]$ code over $\mathbb{F}_{q}$. If $C$ is dual-containing, then there exists a quantum stabilizer code $Q$ with parameters $n,2k-n,\geq d]]_{q}$ that is pure to $d$. Moreover, if the minimum distance of $C^{\perp}$ exceeds $d$, then the quantum code $Q$ has minimum distance $d$.
In recent years, researchers in coding theory have worked diligently to construct good (e.g., quantum MDS) quantum codes from classical Euclidean, Hermitian, or symplectic self-orthogonal, dual-containing, or self-dual codes Abdukhalikov-Bag-Panario2023; Ball21; Cao24; DinhLeNguyenTansuchat21; FangFu19; GalindoHernando; galindo; GaoSunYanZhao25; JinXing13; LiuLiu21; ShiYueChang18; TianZhu19; WanZhengZhu24. Moreover, the following results (which will be used in the examples in Section 6) present new quantum code constructions utilizing the lengthening, shortening, and dimension reduction of a given stabilizer quantum code.
@@
(Ketkar-Klappenecker-Kumar-Sarvepalli2006, Lemmas 69, 70, 71, and Corollary 73)
  1. 1.
@@
  4. 4.
    Suppose that an $n,q^{k},d]]_{q}$ and an $n^{\prime},q^{k^{\prime}},d^{\prime}]]_{q}$ stabilizer code exist. Then there exists an $n+n^{\prime},q^{k+k^{\prime}},\min\{d,d^{\prime}\}]]_{q}$ stabilizer code.
3 A Construction of ESO/EDC/ESD QC Codes From Their Constituent Codes
In this section, we discuss the main results of this paper where we propose a recursive construction of infinite families $\{\mathcal{C}_{u}\}_{u\geq 1}$ of good QC codes that are self-orthogonal (or dual-containing or self-dual) with respect to the Euclidean inner product. Such constructions of QC codes will be utilized to show the existence of QECCs in Section 6.
3.1 Recursive Construction of Infinite Families of QC Codes
To provide a careful description of our construction, we first describe a general case of the construction and then specify certain conditions on the constituent codes to explicitly describe the code dimension and lower bound for its minimum distance. Let $x^{m}-1$ factor into $2t+s$ irreducible factors as in (7), where $f_{s}=x-1$. equation* x^m-1 = δg_1 g_1^* ⋯g_t g_t^* f_1⋯f_s-1 f_s, where $\delta\in\mathbb{F}_{q}$ is nonzero, $g_{1},g_{1}^{*},...,g_{t},g_{t}^{*}$ are the factors grouped in pairs with their reciprocal polynomials for each $1\leq j\leq t$, and $f_{1},...,f_{s}$ are self-reciprocal polynomials for $1\leq i\leq s$, where $f_{s}=x-1$. Then,
$R^{\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}.$ (25)
equation* R^ℓ_q,m ≅(Fq[x]⟨g1Fq[x]⟨g1* )^ℓ⊕⋯⊕( Fq[x]⟨gtFq[x]⟨gt* )^ℓ⊕(Fq[x]⟨f1 )^ℓ⊕⋯⊕(Fq[x]⟨fs-1 )^ℓ⊕(Fq[x]⟨x-1 ⟩)^ℓ. Let $\mathcal{C}_{1}\subseteq\mathbb{F}_{q}^{m\ell}$ be an $\ell$-QC code whose first $2t$ constituent codes are direct sums of arbitrary codes and their duals, i.e.,
$\bigoplus_{j=1}^{t}\left(C^{\prime}_{1,j}\oplus{C^{\prime}}_{1,j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{\ell}$
@@ 3.1 Recursive Construction of Infinite Families of QC Codes
$\left(\bigoplus_{i=1}^{s-1}C_{1,i}\right)\oplus C_{1,s}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{\ell}$
where the codes $C_{1,i}$ are HSO codes for $1\leq i\leq s-1$ and $C_{1,s}$ is ESO (see Remark 2.9). Thus,
$\mathcal{C}_{1}\cong\Bigg(\bigoplus_{j=1}^{t}\left(C^{\prime}_{1,j}\oplus{C^{\prime}}_{1,j}^{\perp_{E}}\right)\Bigg)\oplus\left(\bigoplus_{i=1}^{s-1}C_{1,i}\right)\oplus C_{1,s}.$ (26)
As $C_{1,s}\subset\mathbb{F}_{q}^{\ell}$, we can assume without much loss of generality that the it is the constituent code of $\mathcal{C}_{1}$ with the smallest minimum distance. Moreover, as the $C_{1,i}$ are HSO for each $1\leq i\leq s-1$ and $C_{1,s}$ is ESO, Theorem 2.8 tells us that $\mathcal{C}_{1}$ is an ESO $\ell$-QC code of length $m\ell$. Now using $\mathcal{C}_{1}$, we construct another QC code $\mathcal{C}_{2}$ as follows: Let
$R^{m\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}.$
Let $\mathcal{C}_{2}\subseteq\mathbb{F}_{q}^{m^{2}\ell}$ be the $m\ell$-QC code whose first $2t$ constituent codes are the following direct sums of arbitrary codes and their duals,
$\bigoplus_{j=1}^{t}\left(C^{\prime}_{2,j}\oplus{C^{\prime}}_{2,j}^{\perp_{E}}\right)\subset\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m\ell},$
@@
$\left(\bigoplus_{i=1}^{s-1}C_{2,i}\right)\oplus\mathcal{C}_{1}\subset\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}$
@@
where the $C_{2,i}$ are HSO codes for $1\leq i\leq s-1$ and $\mathcal{C}_{1}\subset\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m\ell}\cong\mathbb{F}_{q}^{m\ell}$ is an ESO constituent code of $\mathcal{C}_{2}$. Thus,
@@
$\mathcal{C}_{2}\cong\Bigg(\bigoplus_{j=1}^{t}\left(C^{\prime}_{2,j}\oplus{C^{\prime}}_{2,j}^{\perp_{E}}\right)\Bigg)\oplus\left(\bigoplus_{i=1}^{s-1}\;C_{2,i}\right)\oplus\mathcal{C}_{1}$
@@
Again by Theorem 2.8, $\mathcal{C}_{2}$ is an ESO $m\ell$-QC code of length $m^{2}\ell$. To extend this recursive construction, we have that any $u\geq 2$,
@@
$R^{m^{u-1}\ell}_{q,m}\cong\bigoplus_{j=1}^{t}\left(\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}\oplus\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}\right)^{m^{u-1}\ell}\oplus\left(\bigoplus_{i=1}^{s-1}\left(\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}\right)^{m^{u-1}\ell}\right)\oplus\left(\frac{\mathbb{F}_{q}[x]}{\langle x-1\rangle}\right)^{m^{u-1}\ell}.$
@@
and define an $m^{u-1}\ell$-QC code $\mathcal{C}_{u}$ in $\mathbb{F}_{q}^{m^{u}\ell}$ equation C_u ≅(C’_u1⊕C’_u1^⟂) ⊕(C’_u2⊕C’_u2^⟂) ⊕⋯⊕(C’_u,t⊕C’_u,t^⟂) ⊕C_u1 ⊕⋯⊕C_u,s-1 ⊕C_u-1, whose first $2t$ arbitrary constituent codes are